# MiMo-V2-Flash Technical Report

### v1

@@ 4.6 RL Infrastructures
Our RL (and MOPD) infrastructure uses SGLang (zheng2024sglang) as the inference engine and Megatron-LM (shoeybi2019megatron) as the training engine. We adopt FP8 for both training and inference. To enable stable, efficient, and flexible RL training, we implement three extended modules: Rollout Routing Replay (ma2025stabilizing) (Sec 4.6.1), Data Scheduler (Sec 4.6.2), and Tool Manager combined with Toolbox (Sec 4.6.3).
@@ 4.6.1 Stablized Training via Rollout Routing Replay (R3)
MoE models suffer from inconsistent expert routing across rollout and training due to numerical precision issues (he2025nondeterminism; yao2025offpolicy). We propose Rollout Routing Replay (R3) (ma2025stabilizing) to train RL using the same routed experts from rollout, making its overhead negligible through optimized data types and communication overlapping. For multi-turn agent training, we employ a request-level prefix cache during rollout. This cache stores KVCache and MoE routed experts from prior turns, allowing them to be reused for subsequent generation steps of the same request. Unlike the commonly-used radix cache in current inference engines, our request-level prefix cache avoids re-prefilling or inter-request output cache sharing, ensuring sampling consistency for routed experts.
@@ 4.6.2 Data Scheduler
For MiMo-V2-Flash, we extend the Seamless Rollout Engine (xia2025mimo) and implement a Data Scheduler to seamlessly schedule fine-grained sequences instead of micro-batches, addressing GPU idleness in distributed MoE training. In dynamic sampling, as sequences return for reward computation, we reference historical pass rates and, if necessary, assign new prompts to GPUs with load balancing. We integrate partial rollout (team2025kimi; fu2025areal) to partition overlong trajectories across steps, while limiting staleness and the proportion of partial samples in each batch. By employing staleness-aware truncated importance sampling for partial rollout, we significantly accelerate RL training without sacrificing model quality.
@@ 4.6.3 Toolbox and Tool Manager
We implement Toolbox and Tool Manager to tackle global resource contention and local inefficiency in RL agent training. These modules leverage Ray (moritz2018ray) for efficient scheduling. Toolbox acts as the centralized resource allocator, enforcing resource quota and QPS limits for tools across concurrent tasks. It adopts fault-tolerant Ray actor pools, which eliminate cold-start delays. Integrated with the rollout engine, Tool Manager coordinates with Toolbox to accelerate training through environment pre-warming and sequence-level asynchronous reward computation. It maintains training stability through timeout recovery and real-time monitoring. By disaggregating the tool management and rollout workflow, Toolbox isolates task-specific logic from system-wide policies, enabling modular extensibility without compromising stability.

### v2

@@ 4.6 RL Infrastructures
Our RL (and MOPD) infrastructure uses SGLang [zheng2024sglang] as the inference engine and Megatron-LM [shoeybi2019megatron] as the training engine. We adopt FP8 for both training and inference. To enable stable, efficient, and flexible RL training, we implement three extended modules: Rollout Routing Replay [ma2025stabilizing] (Sec 4.6.1), Data Scheduler (Sec 4.6.2), and Tool Manager combined with Toolbox (Sec 4.6.3).
@@ 4.6.1 Stablized Training via Rollout Routing Replay (R3)
MoE models suffer from inconsistent expert routing across rollout and training due to numerical precision issues [he2025nondeterminism, yao2025offpolicy]. We propose Rollout Routing Replay (R3) [ma2025stabilizing] to train RL using the same routed experts from rollout, making its overhead negligible through optimized data types and communication overlapping. For multi-turn agent training, we employ a request-level prefix cache during rollout. This cache stores KVCache and MoE routed experts from prior turns, allowing them to be reused for subsequent generation steps of the same request. Unlike the commonly-used radix cache in current inference engines, our request-level prefix cache avoids re-prefilling or inter-request output cache sharing, ensuring sampling consistency for routed experts.
@@ 4.6.2 Data Scheduler
For MiMo-V2-Flash, we extend the Seamless Rollout Engine [xia2025mimo] and implement a Data Scheduler to seamlessly schedule fine-grained sequences instead of micro-batches, addressing GPU idleness in distributed MoE training. In dynamic sampling, as sequences return for reward computation, we reference historical pass rates and, if necessary, assign new prompts to GPUs with load balancing. We integrate partial rollout [team2025kimi, fu2025areal] to partition overlong trajectories across steps, while limiting staleness and the proportion of partial samples in each batch. By employing staleness-aware truncated importance sampling for partial rollout, we significantly accelerate RL training without sacrificing model quality.
@@ 4.6.3 Toolbox and Tool Manager
We implement Toolbox and Tool Manager to tackle global resource contention and local inefficiency in RL agent training. These modules leverage Ray [moritz2018ray] for efficient scheduling. Toolbox acts as the centralized resource allocator, enforcing resource quota and QPS limits for tools across concurrent tasks. It adopts fault-tolerant Ray actor pools, which eliminate cold-start delays. Integrated with the rollout engine, Tool Manager coordinates with Toolbox to accelerate training through environment pre-warming and sequence-level asynchronous reward computation. It maintains training stability through timeout recovery and real-time monitoring. By disaggregating the tool management and rollout workflow, Toolbox isolates task-specific logic from system-wide policies, enabling modular extensibility without compromising stability.